# Understanding the Issue Types in Open Source Blockchain-based Software Projects with the Transformer-based BERTopic

### v1

@@ V-A Limitations
A limitation of this study is its exclusive focus on open-source blockchain projects, which may not fully capture the breadth of development practices and challenges across the broader blockchain ecosystem. Another limitation stems from relying solely on issue titles for topic modeling, due to the common presence of auto-generated or repetitive templates in issue bodies. However, many issues include rich, developer-written content that could provide deeper insights.
V-B Future Work
@@
Future research can expand on this study by comparing blockchain-based projects with non-blockchain projects to highlight the unique challenges of blockchain technology. Exploring distinct blockchain domains, such as cryptocurrency, healthcare, finance, supply chain, and smart contracts, would provide a more comprehensive view of industry-specific issues. Investigating closed-source blockchain projects could further enrich our understanding by contrasting them with open-source environments. Finally, the quality of generated topics could be improved by selectively incorporating informative issue bodies while using automated filtering to exclude templated or auto-generated content.
@@ VI Conclusion
This study investigates the challenges in blockchain-based software applications by mining and analyzing roughly half a million GitHub issues from more than a thousand open-source projects. Using topic modeling with BERTopic followed by manual annotation, 49 distinct topics were identified, which were further organized into a hierarchical structure comprising 11 major subcategories. General software issues slightly outnumber blockchain-specific ones, with frequent concerns including logging, wallet management, and UI enhancement. Issue resolution times varied, with wallet-related issues taking longer. While issue activity peaked around 2016 and declined after 2022, wallet issues remained consistently significant. The results show that blockchain development shares challenges with general software engineering but also faces unique issues that require dedicated solutions and practices.

### v2

@@ V-A Limitations & Future Work
This study has several limitations that suggest directions for future research. First, its focus on open-source projects may not fully capture the breadth of development practices and challenges across the broader blockchain ecosystem. Another limitation arises from relying solely on issue titles for topic modeling, as they often contain auto-generated or template text in issue bodies. However, many issues include rich, developer-written content that could provide deeper insights.
@@
Future research could address these limitations by incorporating closed-source blockchain projects to compare practices across different development environments. Expanding the scope to include non-blockchain projects would help highlight challenges unique to blockchain technology. Exploring distinct blockchain domains, such as cryptocurrency, healthcare, finance, supply chain, and smart contracts, would provide a more comprehensive view of industry-specific issues. Finally, the quality of generated topics could be improved by incorporating informative issue bodies while using automated filtering to exclude templated or auto-generated content.
@@ VI Conclusion
This study investigates the challenges in blockchain-based software applications by mining and analyzing roughly half a million GitHub issues from more than a thousand open-source projects. Using topic modeling with BERTopic followed by manual annotation, 49 distinct topics were identified, which were further organized into a hierarchical structure comprising 11 major subcategories. General software issues slightly outnumber blockchain-specific ones, with frequent concerns including logging, wallet management, and UI enhancement. Issue resolution times varied, with wallet-related issues taking longer. While issue activity peaked around 2016 and started declining after 2022, wallet issues remained consistently significant. The results show that blockchain development shares challenges with general software engineering but also faces unique issues that require dedicated solutions and practices.